What are the values of B, and Br? [23] =]。 A B = B. - 3) B = % -- CC B = - B. -- {D} B =10; i = 0
Answer #1

77 B, - 3 L B₂ da. 9B, +7 By 22 -> ☺ B + 3 By =1 + 9 B, 4 7 B2=2 9B,+ 27 B₂ = 9 -20 B2 = -7 -> B2- i form @ B,-,- 3B, -/- 2 2

Hope it will help you.